5x-3y=16
4x+5y=-2

Answer by stanbon(75887)

5x-3y=16
4x+5y=-2
----------------------
Modify:
25x - 15y = 80
12x + 15y = -6
-----
Add and solve for "x":
37x = 74
---
x = 2
----
Solve for "y":
4x + 5y = -2
8 + 5y = -2
5y = -10
y = -2
=======================
Cheers,
Stan H.
=======================